Finish the sequence 3. 6. 18 ___
which of these numbers is correct
84
60
72
30

Answer by math_tutor2020(3844)


This sequence could be geometric.
To find out, divide adjacent terms
term2/term1 = 6/3 = 2
term3/term2 = 18/6 = 3
The ratios we got were 2 and 3.
We do not get the same ratio each time, so the sequence isn't geometric.

But the pattern could be: multiply by 2, multiply by 3, multiply by 4, etc
If so, then
18*4 = 72
could be the next term.

Unfortunately sequence problems like this are very flawed.
Check out this similar problem
https://www.algebra.com/algebra/homework/Sequences-and-series/Sequences-and-series.faq.question.1195799.html
On that link I try to find the next few terms in the sequence 1,2,4,...
It turns out there are at least 3 different possible answers for that question. There may be infinitely many answers.
Without knowing the actual rule to "3, 6, 18, ..." there's no way of knowing 100% what the next term could be.